In a young double slit experiment the separation is 0.05cm and the distance between the double slit and screen is 200cm when blue light is used , the distance of the interference pattern is 0.13cm


1
Expert's answer
2021-03-16T09:13:52-0400

Given,

Seperation d=0.05 cm


D=200 cm

x=0.13 ,

n=1


For bright fringes, "x=\dfrac{nλD}{d} or λ=\dfrac{xd}{nD}"



Substituting the values, we get


"\lambda=\dfrac{0.13\times 0.05}{ 1\times 200}=\dfrac{65\times 10^{-4}}{200}=0.325\times 10^{-4}cm"